OpenStudy (anonymous):

Please Help!!!!!!! Solve each system using elimination. Tell whether each system has no solution, infinitely many solutions, or one solution. 1. x+2y=23 5x+10y=55 2. 5x+4y=-83 3x-3y=-12 3. 4x+y=21 -2x+6y=9

OpenStudy (anonymous):

are you familiar with elimination system?

OpenStudy (anonymous):

Yes I just have a lot of work to do and don't have time to go over it step-by-step. My deadline is tonight.

OpenStudy (usukidoll):

I don't think I could use elimination on these.. I can't find a common number... subsitution would be better

OpenStudy (anonymous):

you can @UsukiDoll

OpenStudy (usukidoll):

well ok maybe but isn't that for elimination you need like a 5x on the first equation and -5x on the second... that's just an example btw

OpenStudy (anonymous):

Well can any of you help me???????

OpenStudy (anonymous):

x+2y = 23 -- 5x+10y = 55 multiply the first equation by 5 then subtract it the 2nd equation, see what you get 5x+10y = 115 5x+10y = 55 now subtract them

OpenStudy (anonymous):

40

OpenStudy (anonymous):

no, subtract the like terms 5x+10y = 115 - 5x+10y = 55 -------------- 0 + 0 = 60 --- so this would have no solution

OpenStudy (anonymous):

Ok I get it...

OpenStudy (anonymous):

both the variables cancels out when ever you are trying to eliminate them, which leaves them with no solution

OpenStudy (anonymous):

for the second one, think of which variables you want to eliminate first then find there common number then either subtract of add depending on their sign

OpenStudy (anonymous):

. 5x+4y=-83 3x-3y=-12 Eliminate 4y and 3y and 5x and 3x I think...

OpenStudy (anonymous):

you have 5x+4y=-83 3x-3y=-12 so lets say, you want to eliminate the x first so you would find 5 and 3 common number

OpenStudy (anonymous):

So what do u mean by common number?

OpenStudy (anonymous):

5 and 3 had common number of 15 because if you multiply 5 and 3 you would get 15

OpenStudy (anonymous):

Ohh yaaa...

OpenStudy (anonymous):

so you would multiply 5x+4y=-83--- multiply by 3 3x-3y=-12--- multiply by 5

OpenStudy (anonymous):

S0 4y and 3y common number is 12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So basically 5x+4y=-83 and 3x-3y=-12 has 1 solution?

OpenStudy (anonymous):

I graphed it and got only 1 solution...

OpenStudy (anonymous):

yeah, but dont you want to know how to solve it 5x+4y=-83--- multiply by 3 3x-3y=-12--- multiply by 5 so 15x+12y = -249 15x-15y = -60 --- subtract them --------------- 0 - 27y = -189 -- divide both sides by 27 y = -7

OpenStudy (anonymous):

Yes I get that and now I need the 3rd one...

OpenStudy (anonymous):

do the same thing we did for the 2nd one first pick a variable to eliminate and then find their common number then subtract or add the 2 equations depending on their signs

OpenStudy (anonymous):

4x+y=21 -2x+6y=9 Like terms- x= -8 and for y it is just 6? as there i no value for y in the 1st equation

OpenStudy (anonymous):

the coefficient for the first y is just 1

OpenStudy (anonymous):

Ok so the y like term is 6?

OpenStudy (anonymous):

So the like terms are -8 and 6?

OpenStudy (anonymous):

it can be 4 and 6, because when you multiply 2 by 2, you would get 4, so you dont really need to do anything to the first equation to eliminate the x

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so 4x+y=21 -2x+6y=9--- multiply by 2 4x+y=21 -4x+12y=18 now since the second 4 is negative, add them ------------- 0 +13y = 39 divide both by 13 y = 3

OpenStudy (anonymous):

do the same process to find the x eliminate the y then solve for x

OpenStudy (anonymous):

I get it now....thx!

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

So is x 3?

OpenStudy (anonymous):

no it should be 4x+3 = 21 4x = 18 x=4.5

OpenStudy (anonymous):

Oh yes I messed up in the multiplying part :P I am just soo overwhelmed with the work I have...

OpenStudy (anonymous):

So 1 solution right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ok thx!!